OpenStudy (coconut):

What is r when (2x3 + 4x2 – 32x + 18) ÷ (x + 3)? I believe r means remainder

OpenStudy (paxpolaris):

\[\large r=2\left( -3 \right)^3+4\left( -3 \right)^2-32\left( -3 \right)+18\]

OpenStudy (paxpolaris):

that is if you know the remainder theorem

OpenStudy (coconut):

is the answer 96 ?

OpenStudy (anonymous):

yes.